OpenStudy (anonymous):

Find all possible solutions for the triangle, A=50º, a=15, b=12. SOmeone please explain how to do this type of problem

OpenStudy (anonymous):

law of sines yes? first step is to draw a reasonable triangle.

OpenStudy (anonymous):

yes, ok done

OpenStudy (anonymous):

you should have two triangles i think. do you?

OpenStudy (anonymous):

why 2 triangles?

OpenStudy (anonymous):

hold on one sec i will see if i can draw it for you somehow

OpenStudy (anonymous):

A = 50o a = 12 b = 15 could look like this one exactly yes?

OpenStudy (anonymous):

o, i just drew a simple triangle like this

OpenStudy (anonymous):

yeah a little hard for me to see, but the image i sent you, that i just found by googling law of sines two triangles" looks pretty much exactly like what we are going to solve. so lets use that one

OpenStudy (anonymous):

put A = 50, a = 12 b = 15 you can see that you can pivot the 12 side to make two triangles

OpenStudy (anonymous):

first we solve for the big triangle like the one you drew.

OpenStudy (anonymous):

law of sines gives \[\frac{\sin(B)}{b}=\frac{\sin(A)}{a}\] \[\sin(B)=\frac{b\sin(A)}{a}\] \[\sin(B)=\frac{15\sin(50)}{12}\]

OpenStudy (anonymous):

we want B not the sine of B so take the inverse sine of that mess here i did it in one step http://www.wolframalpha.com/input/?i=arcsine%2815+sin%2850+degrees%29%2F12%29

OpenStudy (anonymous):

looks like B = 73.25

OpenStudy (anonymous):

angle C is now easy since they have to add up to 180

OpenStudy (anonymous):

180-50-75.25= 56.75

OpenStudy (anonymous):

so we have A= 50 a = 12 B = 73.25 b = 15 C = 56.75

OpenStudy (anonymous):

and we can find c by using the law of sines again

OpenStudy (anonymous):

and then you just do the law of cosine right?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

law of cosines is a pain. only use it when you cannot use the law of sines

OpenStudy (anonymous):

oo, ok

OpenStudy (anonymous):

for c we just write \[\frac{c}{\sin(C)}=\frac{a}{\sin(A)}\] and since we know 3 out of these 4 numbers we are good to go

OpenStudy (anonymous):

\[c=\frac{12\sin(56.75)}{\sin(50)}\]

OpenStudy (anonymous):

i think you used the length of b instead of a

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

ya. a=15

OpenStudy (anonymous):

damn damn damn

OpenStudy (anonymous):

then first answer was wrong too!

OpenStudy (anonymous):

i got c=16.38

OpenStudy (anonymous):

o darn

OpenStudy (anonymous):

ok we do it quickly now

OpenStudy (anonymous):

i think i can figure out how to solve it now though. thank you!

OpenStudy (anonymous):

\[\sin(B)=\frac{12\sin(50)}{15}\] \[B=\sin^{-1}(\frac{12\sin(50)}{15})\]

OpenStudy (anonymous):

you good from here? this time there is only one triangle

OpenStudy (anonymous):

yup, thanks!